OpenStudy (gabylovesyou):

@johnweldon1993 John and Tim are looking at the equation the square root of the quantity of 3 times x minus 4 equals square root of x . John says that the solution is extraneous. Tim says the solution is non-extraneous. Is John correct? Is Tim correct? Are they both correct? Justify your response by solving this equation, explaining each step with complete sentences.

OpenStudy (gabylovesyou):

@amistre64

OpenStudy (anonymous):

sqrt(3x+4) = sqrt(x),

OpenStudy (gabylovesyou):

yes @kennyboy13

OpenStudy (anonymous):

got it? :)

OpenStudy (gabylovesyou):

how do u solve it

OpenStudy (anonymous):

So, we have that 3x+4 − − − − − √ =x √ If we square both sides, we have 3x+4=x Solving for x, we get x=-2. Substituting into the original equation, we get −2 − − − √ =−2 − − − √ Since the solution gives us a nonreal number, we conclude that the solution is extraneous.